OpenStudy (anonymous):

The perimeter of a rectangle is 64 units. Can the length x of the rectangle can be 20 units when its width y is 11 units? No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 64. No, the rectangle cannot have x = 20 and y = 11 because x + y ≠ 32. Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 64. Yes, the rectangle can have x = 20 and y = 11 because x + y is less than 32.

OpenStudy (anonymous):

@johnweldon1993 help please^.^ sorry...

OpenStudy (johnweldon1993):

Well you know that the perimeter would be 2L + 2W... is L = 20 and W = 11...that would be 2(20) + 2(11) = 40 + 22 = 62 So no it cannot...x + y does not equal 32...

OpenStudy (anonymous):

So B

OpenStudy (johnweldon1993):

Never apologize hun That's correct :)